OpenStudy (anonymous):

What's the answer for - (x - 1) = -4x - 2 ?

OpenStudy (solomonzelman):

\[- (x - 1) = -4x - 2\]\[- x + 1 = -4x - 2\]\[- x +4x = - 2 -1\]\[3x=-3\]\[x=-1\]

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

It is -1, like barron16 said